Suppose the value of a car depreciates linearly for 15 years after it is bought. Also, suppose the value of the new car is $25,000 and the value 10 years after it is bought is $9,000. Let y be the value of the car x years after it is bought. Find a linear equation relating x and y. Write in y = mx + b form. I came up with y=1,600 + 23,400. I don't feel this is correct???
Answer by stanbon(75887)

Suppose the value of a car depreciates linearly for 15 years after it is bought. Also, suppose the value of the new car is $25,000 and the value 10 years after it is bought is $9,000. Let y be the value of the car x years after it is bought. Find a linear equation relating x and y. Write in y = mx + b form. I came up with y=1,600 + 23,400. I don't feel this is correct???
----
You have 2 points relating time and value: (0,25000) and (10,9000)
----
slope = (9000-25000)/(10-0) = 1600
------
intercept is f(0) = 25000
------------
Equation:
value = 1600(years) + 25000
v = 1600x + 25000
=======================
Cheers,
Stan H.
=======================